# The prescribed metric on the boundary of convex subsets of anti-de Sitter space with a quasi-circle as ideal boundary

### v1

@@ 2. Anti-de Sitter space
The anti-de Sitter space $\mathbb{ADS}^{n,1}$ is the Lorentzian analogue of the hyperbolic space $\mathbb{H}^{n+1}$, in the sense that any Lorentzian manifold of constant Gaussian curvature equal to $-1$ is locally modeled on $\mathbb{ADS}^{n,1}$. In this section we will introduce the models for the three dimensional anti-de Sitter space $\mathbb{ADS}^{2,1}$ that we will use later thought this paper. For more details we refer the reader to [BS20] and[Mes07].
@@ 2.1. Hyperbloid model
Let $q_{2,2}$ be the quadratic form defined on $\mathbb{R}^{4}$ by the formula
@@
We define the space $\mathbb{H}^{2,1}$ to be
@@
Then $q_{2,2}$ induce a scalar product on each tangent space of $\mathbb{H}^{2,1}$ that has signature $(2,1)$, this makes $\mathbb{H}^{2,1}$ a Lorentian manifold. We refer to [BS20] and[Mes07] to see why $\mathbb{H}^{2,1}$ has a constant Gaussian curvature equal to $-1$.
@@ 2.2. The projective model of $\mathbb{ADS}^{2,1}$
We introduce $\mathbb{ADS}^{2,1}$, the projective model of the anti-de Sitter space, to be
$\mathbb{ADS}^{2,1}:=\mathbb{H}^{2,1}/\left\{\pm\right\}$
or equivalently
@@
We refer to [BS20, Section 2.2] to see why $\mathbb{ADS}^{2,1}$ has constant sectional curvature equal to $-1$.
@@ 2.3. The Lie group model
There is an explicite identification between $\partial_{\infty}\mathbb{ADS}^{2,1}$ and $\mathbb{RP}^{1}\times\mathbb{RP}^{1}$ via the following map
@@
$\left[M\right]$ $\to(Im(M),Ker(M))$
Note that group $\rm{PSL}(2,\mathbb{R})\times\rm{PSL}(2,\mathbb{R})$ acts on $\rm{PSL}(2,\mathbb{R})$ by left and right composition, that is
$(A,B).X=AXB^{-1}$
@@
In general, in any Lorentzian manifold $(M,q)$, we say that a vector $v\in T_{p}M$ is
@@
The geodesics in $\mathbb{ADS}^{2,1}$ are obtained by the intersection of planes of $\mathbb{R}^{2,2}$ that go thought the origin with $\mathbb{H}^{2,1}$. We say that a geodesic $\alpha$ is
@@ 2.4. Totally geodesic planes
A plane $P$ in $\mathbb{ADS}^{2,1}$ is defined as the intersection of $W$, a three dimensional vector space in $\mathbb{R}^{2,2}$, with $\mathbb{H}^{2,1}$. That is a plane $P$ in $\mathbb{ADS}^{2,1}$ is defined as
$P:=W\cap\mathbb{H}^{2,1}$
We say that a plane $P$ is
@@ 2.6. Acausal curves
We will define the domain of dependence of a quasi-circle $C$. We say that a curve in $\mathbb{ADS}^{2,1}$ is causal if its tangent vector at any point is time-like or light-like
@@
Let $C\subset\partial_{\infty}\mathbb{ADS}^{2,1}$ be a quasi-circle. We define $D(C)$, the domain of dependence of $C$, to be
$D(C):=\left\{p\in\mathbb{ADS}^{2,1},\textit{such that every inextensible no causal curve through $p$ meets $C$}\right\}$
@@
We also define the following
@@
If $C$ is not the graph of a projective map then $\partial\rm{CH}(C)\cap\mathbb{ADS}^{2,1}=\partial^{+}\rm{CH}(C)\bigsqcup\partial^{-}\rm{CH}(C)$, where each of ±$\partial^{\pm}CH(C)$ is a topological disc which is a pleated surface.
@@ 2.7. The width of the convex hull of a meridian
In the next proposition we will use the notion of anti-de Sitter rhombus, for a definition we refer to [BDMS21, Example 6.7].
@@
[BS10] Let $C\subset\partial_{\infty}\mathbb{ADS}^{2,1}$ be an acausal curve. Then $C$ is a quasicircle if and only if $w(C)<\frac{\pi}{2}$. Further, if $C_{n}$ is a sequence of quasicircles whose optimal quasisymmetric constant diverges to infinity, then there exist isometries $\phi_{n}\in\text{Isom}_{\mathbb{ADS}^{2,1}}$ so that $\phi_{n}(C_{n})$ converges to the rhombus $C^{*}$, so that in particular $w(C_{n})\to\frac{\pi}{2}$.
@@ 2.8. The left and right projections
Let $S$ be a surface and let $(M,g)$ be a Lorentzian manifold, a $C^{1}$ immersion $\sigma:S\to M$ is said to be spacelike if the pull back metric $\sigma^{*}(g)$ is a Riemannian metric. A spacelike surface is the image of a $C^{1}$ spacelike embedding. Moreover an immersed surface is convex if its principal curvatures have the same sign (the sign can be positive or negative depending on the orientation).
@@
Recall (see [BS20, Proposition 3.5.2] for example) that any time-like geodesic $L$ in $\mathbb{ADS}^{3}$ is of the form $L=\left\{A\in\rm{PSL}(2,\mathbb{C}),Ax=x^{\prime}\right\}$ for some $x,x^{\prime}\in\mathbb{H}^{2}$. Since at each point $p$ of the surface $S$ there is a unique time-like geoedsic $L_{p}$ that goes thought $p$ and orthogonal to $T_{p}S$, we have a well defined map
@@
$\Pi^{*}_{l}(h_{-1})(v,w)$ $=I\left((E+J_{I}B)v,(E+J_{I}B)w\right)$
$\Pi^{*}_{r}(h_{-1})(v,w)$ $=I\left((E-J_{I}B)v,(E-J_{I}B)w\right)$
@@ 3. Globally hyperbolic convex subsets
Next we define the notion of a globally hyperbolic convex subset
@@
  • $\partial\Omega\cap\mathbb{ADS}^{2,1}$ is the disjoint union of two smooth spacelike disks ±$\partial^{\pm}\Omega$.
  • The induced metric on each of ±$\partial^{\pm}\Omega$ has curvature in the interval $(-\frac{1}{\epsilon},-1-\epsilon)$
@@ 4.1. Definition of gluing maps
Under some conditions on the induced metrics on the boundary of $\Omega$, the principal curvatures of ±$\partial^{\pm}\Omega$ are always bounded.
@@
[BDMS21, Lemma 7.9] Let $S_{n}$ be a sequence of properly embedded convex spacelike disks spanning a sequence of $k$-quasicircles $C_{n}$. If $C_{n}$ converges to the union $\{p\}\times\mathbb{RP}^{1}\cup\mathbb{RP}^{1}\times\{q\}$ of a line of the left ruling and a line of the right ruling, then $S_{n}$ converges to the lightlike plane with boundary at infinity $\{p\}\times\mathbb{RP}^{1}\cup\mathbb{RP}^{1}\times\{q\}$. If $C_{n}$ converges to a $k$-quasicircle $C$, then, up to a subsequence, $S_{n}$ converges to a locally convex properly embedded surface spanning the curve $C$.
Then we use similar arguments to [BDMS21, Proposition 7.10] and [BDMS21, Lemma 7.11] to show the following proposition
@@
Let $k>1$ and $\epsilon>0$. Let $S\subset\mathbb{ADS}^{2,1}$ be a properly embedded convex spacelike disk spanning a $k$-quasicircle $C$. Assume that $S$ is isometric to $(\mathbb{D},h)$ where $h$ is a conformal complete metric that has curvature in $\left(-\frac{1}{\epsilon},-1-\epsilon\right)$. Assume that there is a sequence $(M_{p})_{p\in\mathbb{N}}$ of positive numbers such that each derivative of $h$ of order $p$ is bounded by $M_{p}$. Then there exists $D$ that depends on $k$, $\epsilon$, and $(M_{p})_{p\in\mathbb{N}}$ such that the principal curvatures of $S$ are in the interval $\left(\frac{1}{D},D\right)$.
@@
We argue by contradiction. Assume the existence of a sequence of properly embedded convex spacelike disks $S_{n}$ spanning $k$-quasicircles $C_{n}$, satisfying the hypothesis of Proposition 4.2, such that there exists a sequence of points $p_{n}\in S_{n}$ in which one of the principal curvatures at $p_{n}$ goes to infinity. Note that the product of the principal curvatures is bounded (because the curvatures belong to $(-\frac{1}{\epsilon},-1-\epsilon)$). If one principal curvature goes to 00, then the other goes to $\infty$. So assume without loss of generality that the largest principal curvature at $p_{n}$ goes to $\infty$.
Up to normalization by isometries, we can assume that $p_{n}$ are equal to a fixed point $p$, and $T_{p_{n}}S_{n}$ are equal to a fixed space-like tangent plane. By Proposition 4.1 and since all $S_{n}$ are tangent to the same space-like tangent plan, the surfaces $S_{n}$ converge in the Hausdorff sense to a locally convex properly embedded surface spanning a $k$-quasi-circle $C$.
@@
Then, up to extracting a subsequence, $h^{\prime}_{n}$ converges smoothly ($C^{\infty}$ on compact subsets) to a metric $h^{\prime}_{\infty}$. By [Sch96, Theorem 5.6], either $\phi_{n}:(\mathbb{D},h^{\prime}_{n})\to S_{n}$ converge smoothly to an isometric embedding $\phi_{\infty}:(\mathbb{D},h_{\infty})\to S_{\infty}$, or there is a complete geodesic $\gamma$ in $(\mathbb{D},h_{\infty})$ which is sent by $\phi_{\infty}$ to a spacelike geodesic $\Gamma\in\mathbb{ADS}^{2,1}$. Moreover, in the latter case, the integral of the mean curvatures of $\phi_{n}$ goes to $\infty$ at any point in a neighborhood of $\gamma$. According to [Sch96, Lemma 5.4], the length of the third fundamental form for $\phi_{n}$ of any geodesic segment transverse to $\gamma$ also goes to $\infty$. This implies that the limit surface $S_{\infty}$ must be contained in the union of the past-directed lightlike half-planes bounded by $\Gamma$. However, this contradicts the fact that the ideal boundary of $S_{\infty}$ is a quasi-circle (Proposition 4.1).
Therefore, the isometries $\phi_{n}:(\mathbb{D},h^{\prime}_{n})\to S_{n}$ must converge, up to extracting a subsequence, to an isometric embedding $\psi_{\infty}:(\mathbb{D},h^{\prime}_{\infty})\to S_{\infty}$. This contradiction confirms that the principal curvatures of $S_{n}$ at $p$ cannot go to $\infty$. ∎
Remark 4.3.
Note that the main reason for assuming that the derivatives of $h$ are bounded is to use [Sch96, Theorem 5.6]. However, in the proof of [Sch96, Theorem 5.6], we need only the convergence of the 2-jets of the metrics. Therefore, we believe that we can relax the condition of boundedness of the derivatives of $h$. Instead of saying that the derivatives of $h$ are bounded at any order, we say that the derivatives of $h$ are bounded at some order $k$. We need to be sure that if $h_{n}$ converge in $C^{k}$, then their 2-jets converge. We believe that it might be sufficient to say that the derivatives of $h$ are bounded up to order 3.
From that, we will deduce that the gluing maps are well defined. Before that, we need to show the following lemma
@@
Let $\Omega$ be a globally hyperbolic convex subset spanning a $k$-quasi-circle at infinity, and let the induced metric on its spacelike boundaries have curvatures in $(-\frac{1}{\epsilon},-1-\epsilon)$. Let $V^{\pm}:(\mathbb{D},h^{\pm})\to\partial^{\pm}\Omega$ be isometries. Assume that each derivative of $h^{\pm}$ at order $p$ is bounded by $M_{p}$ on $\mathbb{D}$. Let $\Pi^{\pm}_{l}$, $\Pi^{\pm}_{r}$ be the left and right projections. Then there exists $A>0$ depending only on $\epsilon$, $k$, and $(M_{p})_{p\in\mathbb{N}}$ such that $\Pi^{\pm}_{l}\circ V^{\pm}$ and $\Pi^{\pm}_{r}\circ V^{\pm}$ are $A$-quasi-isometries.
@@
From Proposition 4.2, there exists $D>0$ depending only on $k$, the quasi-symmetric constant of $\partial_{\infty}\Omega$, $\epsilon$, and $(M_{p})_{p\in\mathbb{N}}$ such that the principal curvatures of ±$\partial^{\pm}\Omega$ are in the interval $\left[\frac{1}{D},D\right]$. By Theorem 2.7, we obtain that the projection maps $\Pi_{l}^{\pm}$, $\Pi_{r}^{\pm}$ are $A$-bilipschitz for some constant $A$ that depends on $D$ (which in turn depends on $\epsilon$, $k$, and $(M_{p})_{p\in\mathbb{N}}$).

### v2

@@ 3. Anti-de Sitter space
The anti-de Sitter space $\mathbb{ADS}^{n,1}$ is the Lorentzian analogue of the hyperbolic space $\mathbb{H}^{n+1}$, in the sense that any Lorentzian manifold of constant sectional curvature equal to 1-1 is locally modeled on $\mathbb{ADS}^{n,1}$. In this section we will introduce the models for the three dimensional anti-de Sitter space $\mathbb{ADS}^{2,1}$ that we will use later thought this paper. For more details we refer the reader to [BS20] and[Mes07].
@@ 3.1. Hyperbloid model
Let $q_{2,2}$ be the quadratic form defined on $\mathbb{R}^{4}$ by the formula :
@@
We define the space $\mathbb{H}^{2,1}$ to be :
@@
Then $q_{2,2}$ induce a scalar product on each tangent space of $\mathbb{H}^{2,1}$ that has signature $(2,1)$, this makes $\mathbb{H}^{2,1}$ a Lorentian manifold. We refer to [BS20] and[Mes07] to see why $\mathbb{H}^{2,1}$ has a constant sectional curvature equal to 1-1.
@@ 3.2. The projective model of $\mathbb{ADS}^{2,1}$
We introduce $\mathbb{ADS}^{2,1}$, the projective model of the anti-de Sitter space, to be :
$\mathbb{ADS}^{2,1}:=\mathbb{H}^{2,1}/\left\{\pm\right\},$
or equivalently :
@@ 3.3. The Lie group model
There is an explicite identification between $\partial_{\infty}\mathbb{ADS}^{2,1}$ and $\mathbb{RP}^{1}\times\mathbb{RP}^{1}$ via the following map :
@@
$\left[M\right]$ $\to(Im(M),Ker(M)).$
Note that group $\rm{PSL}(2,\mathbb{R})\times\rm{PSL}(2,\mathbb{R})$ acts on $\rm{PSL}(2,\mathbb{R})$ by left and right composition, that is :
$(A,B).X=AXB^{-1}.$
@@
In general, in any Lorentzian manifold $(M,q)$, we say that a vector $v\in T_{p}M$ is :
@@
The geodesics in $\mathbb{ADS}^{2,1}$ are obtained by the intersection of planes of $\mathbb{R}^{2,2}$ that go thought the origin with $\mathbb{H}^{2,1}$. We say that a geodesic $\alpha$ is :
@@ 3.4. Totally geodesic planes
A plane $P$ in $\mathbb{ADS}^{2,1}$ is defined as the intersection of $W$, a three dimensional vector space in $\mathbb{R}^{2,2}$, with $\mathbb{H}^{2,1}$. That is a plane $P$ in $\mathbb{ADS}^{2,1}$ is defined as :
$P:=W\cap\mathbb{H}^{2,1}.$
We say that a plane $P$ is :
@@ 3.5. Acausal curves
We will define the domain of dependence of a quasi-circle $C$. We say that a curve in $\mathbb{ADS}^{2,1}$ is causal if its tangent vector at any point is time-like or light-like.
@@
Let $C\subset\partial_{\infty}\mathbb{ADS}^{2,1}$ be a quasi-circle. We define $D(C)$, the domain of dependence of $C$, to be:
[…]
Equivalently, $D(C)$ (see [BB09]) is the unique maximal (in the sense of inclusion) open convex subset whose boundary at infinity is equal to $C$.
@@
We also give the following definition.
@@
If $C$ is not the graph of a projective map then $\partial\rm{CH}(C)\cap\mathbb{ADS}^{2,1}=\partial^{+}\rm{CH}(C)\bigsqcup\partial^{-}\rm{CH}(C)$, where each of $\partial^{\pm}CH(C)$ is a topological disc which is a pleated surface.
@@ 3.6. The width of the convex hull of a meridian
In the next proposition, we will use the notion of a rhombus in $\partial_{\infty}\mathbb{ADS}^{2,1}$, so we define it below before stating the proposition. See also [BDMS21, Example 6.7].
Definition 3.5.
Recall that $\partial_{\infty}\mathbb{ADS}^{2,1}$ is identified with $\mathbb{RP}^{1}\times\mathbb{RP}^{1}$. We call the lines of the form $\{p\}\times\mathbb{RP}^{1}$ (resp $\mathbb{RP}^{1}\times\{p\}$) the right (resp left) ruling lines, where $p\in\mathbb{RP}^{1}$.
Let $a,b,a^{\prime},b^{\prime}\in\mathbb{RP}^{1}$ such that $a<a^{\prime}$ and $b<b^{\prime}$ in the cyclic order of $\mathbb{RP}^{1}$. We denote by $C_{\diamond}$ the piecewise linear curve in $\partial_{\infty}\mathbb{ADS}^{2,1}$ consisting of the horizontal segments (that is, segments lying on a right ruling line) connecting $(a,b)$ to $(a,b^{\prime})$ and $(a^{\prime},b)$ to $(a^{\prime},b^{\prime})$, and the vertical segments (that is, segments lying on a left ruling line) connecting $(a,b)$ to $(a^{\prime},b)$ and $(a,b^{\prime})$ to $(a^{\prime},b^{\prime})$. We call such a curve a rhombus.
Note that there is a unique rhombus up to isometry (that is, up to the action of isometries of $\mathbb{ADS}^{2,1}$). The convex hull $\rm{CH}(C_{\diamond})$ of a rhombus in $\partial_{\infty}\mathbb{ADS}^{2,1}$, taken inside $\mathbb{ADS}^{2,1}$, is a tetrahedron (see [BDMS21, Figure 3]). Since the spacelike segment (inside $\mathbb{ADS}^{2,1}$) connecting $(a,b)$ to $(a^{\prime},b^{\prime})$ is dual to the spacelike segment connecting $(a^{\prime},b)$ to $(a,b^{\prime})$, it follows that $w(C_{\diamond})=\frac{\pi}{2}$. Therefore, the width of a rhombus curve in $\partial_{\infty}\mathbb{ADS}^{2,1}$ is equal to $\frac{\pi}{2}$.
@@
[BS10] Let $C\subset\partial_{\infty}\mathbb{ADS}^{2,1}$ be an acausal curve. Then $C$ is a quasicircle if and only if $w(C)<\frac{\pi}{2}$. Further, if $C_{n}$ is a sequence of quasicircles whose optimal quasisymmetric constant diverges to infinity, then there exist isometries $\phi_{n}\in Isom({\mathbb{ADS}^{2,1}})$ so that $\phi_{n}(C_{n})$ converges to the rhombus $C^{*}$, so that in particular $w(C_{n})\to\frac{\pi}{2}$.
@@ 3.7. The left and right projections
Let $S$ be a surface and let $(M,g)$ be a Lorentzian manifold, a $C^{1}$ immersion $\sigma:S\to M$ is said to be spacelike if the pull back metric $\sigma^{*}(g)$ is a Riemannian metric. Moreover an immersed surface is convex if its principal curvatures have the same sign (the sign can be positive or negative depending on the orientation).
@@
Recall (see [BS20, Proposition 3.5.2] for example) that any time-like geodesic $L$ in $\mathbb{ADS}^{3}$ is of the form $L=\left\{A\in PSL(2,\mathbb{R}),Ax=x^{\prime}\right\}$ for some unique $x,x^{\prime}\in\mathbb{H}^{2}$. Since at each point $p$ of the surface $S$ there is a unique time-like geoedsic $L_{p}$ that goes thought $p$ and orthogonal to $T_{p}S$, we have a well defined map :
@@
$\Pi^{*}_{l}(h_{-1})(v,w)$ $=I\left((E+J_{I}B)v,(E+J_{I}B)w\right),$
$\Pi^{*}_{r}(h_{-1})(v,w)$ $=I\left((E-J_{I}B)v,(E-J_{I}B)w\right),$
@@ 4. Globally hyperbolic convex subsets
Next we define the notion of a globally hyperbolic convex subset :
@@
  • $\partial\Omega\cap\mathbb{ADS}^{2,1}$ is the disjoint union of two smooth spacelike disks $\partial^{\pm}\Omega$.
  • The induced metric on each of $\partial^{\pm}\Omega$ is complete and has curvature in the interval $(-\frac{1}{\epsilon},-1-\epsilon)$.
@@ 5.1. Definition of gluing maps
Under some conditions on the induced metrics on the boundary of $\Omega$, the principal curvatures of $\partial^{\pm}\Omega$ are always bounded.
Before stating the next lemma and proposition, we draw the reader’s attention to the fact that in the $\rm{PSL}(2,\mathbb{R})$ model of $\mathbb{ADS}^{2,1}$, a lightlike plane $P$ is the orthogonal to an element $\left\lfloor A\right\rfloor\in\rm{PSL}(2,\mathbb{R})$ such that $\det(A)=0$ (see [DS24a, Section 3.5]). Moreover, the boundary at infinity of a lightlike plane is given by $\partial_{\infty}P:=(\operatorname{Im}(A)\times\mathbb{RP}^{1})\cup(\mathbb{RP}^{1}\times\operatorname{Ker}(A))$ (see [DS24a, Lemma 3.5]).
Lemma 5.1.
[BDMS21, Lemma 6.3] Let $C_{n}$ be a sequence of $k$-quasicircles in $\partial_{\infty}\mathbb{ADS}=\mathbb{RP}^{1}\times\mathbb{RP}^{1}$.
Then there exists a subsequence which converges in the Hausdorff topology to either a $k$-quasi circle $C$, or to the union $\left\{p\right\}\times\mathbb{RP}^{1}\cup\mathbb{RP}^{1}\times\left\{q\right\}$ of a line of the left ruling and a line of the right ruling (see Definition 3.6 for definition of left ruling and right ruling lines).
@@
[BDMS21, Lemma 7.9] Let $S_{n}$ be a sequence of properly embedded convex spacelike disks spanning a sequence of $k$-quasicircles $C_{n}$. If $C_{n}$ converges in the Hausdorff topology to the union $\{p\}\times\mathbb{RP}^{1}\cup\mathbb{RP}^{1}\times\{q\}$ of a line of the left ruling and a line of the right ruling, then $S_{n}$ converges in the Hausdorff topology to the lightlike plane with boundary at infinity $\{p\}\times\mathbb{RP}^{1}\cup\mathbb{RP}^{1}\times\{q\}$. If $C_{n}$ converges in the Hausdorff topology to a $k$-quasicircle $C$, then, up to extracting a subsequence, $S_{n}$ converges in the Hausdorff topology to a locally convex properly embedded surface spanning the curve $C$.
In what follows, we need Theorem 5.5 and Theorem 5.6 from [Sch96, Section 5]. They are not stated in their full generality, we have stated only what is needed for the rest of our proof. Refer to the same reference for more definitions and details. In particular, for us, $(M,g_{0})$ is assumed to be $\mathbb{ADS}^{2,1}$, but in [Sch96], it is assumed to be a more general $3$-dimensional Lorentzian manifold.
Theorem 5.3.
[Sch96, Theorem 5.5] Let $f_{n}:\mathbb{D}\rightarrow M$ be a sequence of uniformly elliptic space-like immersions of the disk such that $(f_{n}^{*}(g_{0}))_{n\in\mathbb{N}}$ converges in $C^{\infty}$ to a metric $g_{\infty}$, and such that the sequence of integrals of the mean curvatures $H_{n}$ is bounded by a constant. Suppose also that $(x_{n})_{n\in\mathbb{N}}$ and $(j^{1}f_{n}(x_{n}))_{n\in\mathbb{N}}$ converge. Then there exists a subsequence of $(f_{n})_{n\in\mathbb{N}}$ that converges in $C^{\infty}$ to an isometric immersion $f_{\infty}$.
Later, Schlenker studies what happens when the assumptions of Theorem 5.3 do not hold.
Theorem 5.4.
[Sch96, Theorem 5.6] Let $f_{n}:\mathbb{D}\to M$ be a sequence of uniformly elliptic space-like immersions of the disk, such that $(f_{n}^{*}(g_{0}))_{n\in\mathbb{N}}$ converges in $C^{\infty}$ to a metric $g_{\infty}$. Let $(x_{n})_{n\in\mathbb{N}}$ be a sequence converging to $x_{\infty}$ such that $(j^{1}f_{n}(x_{n}))_{n\in\mathbb{N}}$ converges, but $(f_{n})_{n\in\mathbb{N}}$ does not converge in $C^{\infty}$ in a neighborhood of $x_{\infty}$. Then there exists a subsequence of $(f_{n})_{n\in\mathbb{N}}$, still denoted by $(f_{n})_{n\in\mathbb{N}}$, a maximal geodesic $\gamma$ through $x_{\infty}$ in $(\mathbb{D},g_{\infty})$, and a geodesic segment $\Gamma$ in $M$, such that $(f_{n}|\gamma)$ converges to an isometry onto $\Gamma$.
Then we use similar arguments to [BDMS21, Proposition 7.10] and [BDMS21, Lemma 7.11] to show the following proposition.
Before proceeding with the proof, we briefly recall the notion of the third fundamental form.
Let $S$ be a smooth, space-like embedded surface in $\mathbb{ADS}^{2,1}$ (satisfying the same conditions as in Proposition 5.5). For any $p\in S$, let $n_{S}(p)$ denote the unit normal vector to $S$ at $p$, pointing to the future. Let $\nabla$ denote the Levi-Civita connection of $\mathbb{ADS}^{2,1}$. We define the shape operator of $S$ by $B_{S}:=-\nabla_{\cdot}n_{S}$, and the third fundamental form by $I\hskip-2.84544ptI\hskip-2.84544ptI_{S}(\cdot,\cdot):=I_{S}(B_{S}\cdot,B_{S}\cdot).$ (where $I_{S}$ is the induced metric on $S$, which is also called the first fundamental form).
We also recall that the product of the eigenvalues of $B_{S}$ (that is $\det(B_{S})$) is equal to $K_{S}-1$, where $K_{S}$ is the Gaussian curvature of $S$ (which may vary over the surface). Half of the sum of the eigenvalues of $B_{S}$ (that is $\frac{\operatorname{tr}(B_{S})}{2}$, where $\operatorname{tr}(B_{S})$ is the trace of $B_{S}$) is called the mean curvature of $S$ (which may also vary over the surface) and is usually denoted by $H_{S}$.
@@
Let $k>1$ and $\epsilon>0$. Let $S\subset\mathbb{ADS}^{2,1}$ be a properly embedded convex space-like disk spanning a $k$-quasicircle $C$. Assume that $S$ is isometric to $(\mathbb{D},h)$ where $h$ is a conformal complete metric that has curvature in $\left(-\frac{1}{\epsilon},-1-\epsilon\right)$. Assume that there is a sequence $(M_{p})_{p\in\mathbb{N}}$ of positive numbers such that each derivative of $h$ of order $p$ is bounded by $M_{p}$. Then there exists $D$ that depends on $k$, $\epsilon$, and $(M_{p})_{p\in\mathbb{N}}$ such that the principal curvatures of $S$ are in the interval $\left(\frac{1}{D},D\right)$.
@@
We argue by contradiction. Assume the existence of a sequence of properly embedded convex space-like disks $S_{n}$ spanning $k$-quasicircles $C_{n}$, satisfying the hypothesis of Proposition 5.5, such that there exists a sequence of points $p_{n}\in S_{n}$ in which one of the principal curvatures at $p_{n}$ goes to infinity. Note that the product of the principal curvatures is bounded (because the curvatures belong to $(-\frac{1}{\epsilon},-1-\epsilon)$). If one principal curvature goes to 0, then the other goes to $\infty$. So assume without loss of generality that the largest principal curvature at $p_{n}$ goes to $\infty$.
Recall that, by Lemma 5.1, $C_{n}$ converge (up to extracting a subsequence) in the Hausdorff topology to a quasi-circle or to the union of a left ruling line and a right ruling line. Up to normalization by isometries, we can assume that $p_{n}$ are equal to a fixed point $p$, and $T_{p_{n}}S_{n}$ are equal to a fixed space-like tangent plane. By Proposition 5.2 and since all $S_{n}$ are tangent to the same space-like tangent plan, the surfaces $S_{n}$ converge in the Hausdorff sense to a locally convex properly embedded surface $S_{\infty}$ spanning a $k$-quasi-circle $C$.
@@
Then, up to extracting a subsequence, $h^{\prime}_{n}$ converges smoothly ($C^{\infty}$ on compact subsets) to a metric $h^{\prime}_{\infty}$. By Theorem 5.3 and Theorem 5.4, either $V^{\prime}_{n}=V_{n}\circ g_{n}:(\mathbb{D},h^{\prime}_{n})\to S_{n}$ converge smoothly to an isometric embedding $V^{\prime}_{\infty}:(\mathbb{D},h^{\prime}_{\infty})\to S_{\infty}$, or there is a complete geodesic $\gamma$ in $(\mathbb{D},h^{\prime}_{\infty})$ such that $V^{\prime}_{n}(\gamma)$ converge to a space-like geodesic segment $\Gamma\in\mathbb{ADS}^{2,1}$.
We argue by contradiction that we are not in the case of Theorem 5.4. Suppose this is the case. Then by Theorem 5.4 there is a complete geodesic $\gamma$ in $(\mathbb{D},h^{\prime}_{\infty})$ such that $V^{\prime}_{n}(\gamma)$ converge to a space-like geodesic $\Gamma$. According to [Sch96, Lemma 5.4] (and also [Sch96, Proposition 5.2]) the integral of the mean curvatures of $S_{n}$ diverges to $\infty$ at every point in a neighborhood of $\gamma$. Since the products of the eigenvalues of $B_{S_{n}}$ (where $B_{S_{n}}$ is the shape operator of $S_{n}$) are bounded from above and below, the biggest eigenvalue of $B_{S_{n}}$ diverge to $\infty$ and the smallest eigenvalue converge to 0. Recall that from [Sch96, Lemma 5.4] $\gamma$ is limit of curves going in the direction of the smallest eigenvalue. In particular, the length with respect to the third fundamental form of $S_{n}$ along any geodesic segment transverse to $\gamma$ tends to infinity. By the remark made in [Sch96], in the paragraphs before and after [Sch96, Theorem 5.6], the surfaces $S_{n}$ accumulate on a pleated surface contained in the past light-like planes directed by the geodesic segment $\Gamma$.
On the other hand, by Proposition 5.2, and since the surfaces $S_{n}$ have intrinsic curvatures that are uniformly positive and uniformly bounded, it follows that if $C_{n}$ converges to a $k$-quasi-circle $C$, then the limit surface $S_{\infty}$ of the sequence $S_{n}$ cannot be on the past light-like planes directed by the geodesic segment $\Gamma$. This implies that, if we were not in the case of Theorem 5.3, the quasi-circles $C_{n}$ would have to converge to the union of a left ruling and a right ruling line.
Moreover, since all the surfaces $S_{n}$ are tangent to the same space-like plane at the point $p$, Proposition 5.2 again ensures that their ideal boundaries $C_{n}$ converge to a $k$-quasi-circle $C$. This contradicts the assumption that the integral of the mean curvatures of $V^{\prime}_{n}$ diverges to $\infty$ at every point in a neighborhood of $\gamma$.
We conclude that we cannot be in the case of Theorem 5.4.
Therefore, up to extracting a subsequence, the isometries $V^{\prime}_{n}:(\mathbb{D},h^{\prime}_{n})\to S_{n}$ converge to an isometric embedding $V^{\prime}_{\infty}:(\mathbb{D},h^{\prime}_{\infty})\to S_{\infty}$. This shows that the principal curvatures of $S_{n}$ at $p$ cannot diverge. ∎
From that, we will deduce that the gluing maps are well defined. Before that, we need to show the following lemma.
@@
Let $\Omega$ be a globally hyperbolic convex subset spanning a $k$-quasi-circle at infinity. Assume that the induced metrics on $\partial^{\pm}\Omega$ have curvatures in $\left(-\frac{1}{\epsilon},-1-\epsilon\right)$.
Let $V^{\pm}:(\mathbb{D},h^{\pm})\to\partial^{\pm}\Omega$ be isometries. Assume that each derivative of $h^{\pm}$ of order $p$ is bounded by $M_{p}$ on $\mathbb{D}$.
Let $\Pi^{\pm}_{l}$ and $\Pi^{\pm}_{r}$ be the left and right projections. Then there exists $A>0$, depending only on $\epsilon$, $k$, and $(M_{p})_{p\in\mathbb{N}}$, such that $\Pi^{\pm}_{l}\circ V^{\pm}$ and $\Pi^{\pm}_{r}\circ V^{\pm}$ are $A$-quasi-isometries.
@@
From Proposition 5.5, there exists $D>0$ depending only on $k$, the quasi-symmetric constant of $\partial_{\infty}\Omega$, $\epsilon$, and $(M_{p})_{p\in\mathbb{N}}$ such that the principal curvatures of $\partial^{\pm}\Omega$ are in the interval $\left[\frac{1}{D},D\right]$. By Theorem 3.8, we obtain that the projection maps $\Pi_{l}^{\pm}$, $\Pi_{r}^{\pm}$ are $A$-bilipschitz for some constant $A$ that depends on $D$ (which in turn depends on $\epsilon$, $k$, and $(M_{p})_{p\in\mathbb{N}}$).